### v3

# Bourbaki modules and the module of Jacobian derivations of projective hypersurfaces

Abstract.
Two properties of projective hypersurfaces related to the module of Jacobian derivations, namely being tame and being plus-one generated, are discussed in this paper. Tame hypersurfaces are related to Bourbaki ideals, and free hypersurfaces are the simplest examples of tame hypersurfaces. Plus-one generated hypersurfaces are the non free hypersurfaces which are closest to the free ones, and it is an open question whether all of them are tame.
Key words and phrases:
projective hypersurface; Jacobian derivations; minimal resolution; singular subscheme
2020 Mathematics Subject Classification:
Primary 14J70; Secondary 32S25,13D02
1. Introduction
Let $S=\mathbb{C}[x_{0},x_{1},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables $x_{0},x_{1},\ldots,x_{n}$ with complex coefficients, and let $V:f=0$ be a reduced hypersurface of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{n}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e., the homogeneous ideal in $S$ spanned by the partial derivatives $f_{j}$ of $f$ with respect to $x_{j}$ for $j=0,\ldots,n$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$ or, equivalently, the module of derivations killing $f$, namely
(1.1) $D_{0}(f)=\{\theta\in Der(S)\ :\ \theta(f)=0\}=\{\rho=(a_{0},a_{1},\ldots,a_{n})\in S^{n+1}\ :\ \sum_{0}^{n}a_{j}f_{j}=0\}.$
This module is also denoted in the literature by $AR(f)$ (i.e., all Jacobian relations for $f$) or by ${\rm Syz}(f)$ (i.e., the Jacobian syzygies of $f$). We say that $V:f=0$ is an $m$-syzygy hypersurface if the module $D_{0}(f)$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},\ldots,\rho_{m}$, of degrees $d_{j}=\deg\rho_{j}$ ordered such that
$d_{1}\leq d_{2}\leq\ldots\leq d_{m}.$
In this note we assume that $V$ is not the cone over a hypersurface in $\mathbb{P}^{n-1}$, hence $d_{1}\geq 1$. We call these degrees $(d_{1},\ldots,d_{m})$ the exponents of the hypersurface $V$ and $\rho_{1},\ldots,\rho_{m}$ a minimal set of generators for the module $D_{0}(f)$. The smallest degree $d_{1}$ is sometimes denoted by ${\rm mdr}(f)$ and is called the minimal degree of a Jacobian relation for $f$.
Since the $S$-module $D_{0}(f)$ has rank $n$, it follows that one always has
$m\geq n.$
The hypersurface $V$ is free when $m=n$, since then $D_{0}(f)$ is a free module of rank n, see for instance [11, 25, 27, 28, 30]. Let $\Sigma$ be the singular subscheme of $V$, defined by the Jacobian ideal. Then $s=\dim\Sigma\leq n-2$ since $V$ is reduced, and note that the Hilbert polynomial $H(M(f),t)$ of the graded $S$-module $M(f)$ has the form
(1.2) $H(M(f),t)=\frac{\deg(\Sigma)}{s!}t^{s}+\ldots$
where $\ldots$ stand for lower degree terms in $t$ and $\deg(\Sigma)$ is by definition the degree of the scheme $\Sigma$, or of the ideal $J_{f}$.
Beyond the free hypersurfaces, there is the class of plus-one generated hypersurfaces, for short POG hypersurfaces, defined in the case of hyperplane arrangements by Takuro Abe in [1]. With our notation, choosing an order on the syzygies $\rho_{j}$’s, but without assuming that the exponents $d_{j}$’s are ordered, a POG hypersurface is a hypersurface $V:f=0$ in $\mathbb{P}^{n}$ such that $m=n+1$ and the second order syzygies of $f$, that is the relations among the minimal generators $\rho_{j}$ are generated by just one relation, namely
(1.3) $b_{1}\rho_{1}+\ldots+b_{n}\rho_{n}+\ell\rho_{n+1}=0,$
where $\ell\in S_{1}$ is a linear form. If we want to emphasize that this definition depends on the choice of the generators, we say that $V:f=0$ is a POG hypersurface with respect to
$\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1}).$
Since the set of generators is minimal, it follows that none of the coefficients $b_{j}$ is a non-zero constant, that is either $b_{j}=0$ or $b_{j}\neq 0$ and $\deg b_{j}\geq 1$. The hypersurface $V$ is called strictly plus-one generated if in addition $\ell\neq 0$. Note that in this case, we get $d_{n+1}\geq d_{j}$ for all $j$ such that $b_{j}\neq 0$, even if we do not order the degrees $d_{j}$. When $n=2$, any POG curve is in fact a strictly POG curve, see [15, Corollary 2.2] for a more general result applying to any 3-syzygy curve, but in higher dimensions this nice property may fail, see Example 5.5 and Question 5.6.
A POG hypersurface with respect to $\overline{\rho}$ such that $d_{1}\leq\ldots\leq d_{n}=d_{n+1}$ is called a nearly-free hypersurface, see [11, Definition 5.3] in the case $n=3$.
In this note we first collect in Section 2 some basic facts on the module $D_{0}(f)$ in the case of curves in $\mathbb{P}^{2}$, that is $n=2$. In particular, we list some properties for free and plus-one generated curves. We will try to extend these properties in the fourth section to the case $n>2$.
The fact that the case $n>2$ is much harder to understand comes perhaps from the following technical point. Assume we fix
$\hat{\rho}=(\rho_{1},\rho_{2},\ldots,\rho_{n-1})$
a subset of $(n-1)$ elements in a minimal set of generating syzygies for $f$, without an ordering of the degrees. Then we have a sequence of graded $S$-modules
[…]
where
$u(a_{1},\ldots,a_{n-1})=\sum_{j=1}^{n-1}a_{j}\rho_{j}$
and the morphism $v$ is defined as follows. For any syzygy $\rho\in D_{0}(f)$, consider the $(n+1)\times(n+1)$-matrix $M(\hat{\rho},\rho)$ where the first row is given by
$(x_{0},x_{1},\ldots,x_{n}),$
which are the components of the Euler derivation $E$, the $j$-th row, for $2\leq j\leq n$ is given by the components of the $j-1$-syzygy $\rho_{j-1}$ fixed above, and the $n+1$-st row by the components of $\rho.$ Then it is easy to show that the polynomial $\det M(\hat{\rho},\rho)$ is divisible by $f$ and we set
$v(\rho)=\frac{\det M(\hat{\rho},\rho)}{f}.$
Then the image of $v$ is a (shifted) ideal in $S(d_{1}+\ldots+d_{n-1}-d+1)$, which we denote by $B(f,\hat{\rho})=\operatorname{im}v(-(d_{1}+\ldots+d_{n-1}-d+1))$, and which is called the Bourbaki ideal of the module $D_{0}(f)$, see [13, 14, 22] for the case $n=2$ and [7] for the case $n=3$. Following the terminology introduced in [7], we give the following.
Definition 1.1.
We say that the hypersurface $V:f=0$ is tame with respect to the subset $\hat{\rho}$ if the sequence (1.4) is exact.
Remark 1.2.
In the literature, there is a notion of tame arrangements or tame hypersurfaces, see [29] or [32], and another notion of tame polynomials, see [3] or [23]. Our notion of tameness does not seem to be related to either of these notions. Note also that only for hypersurfaces $V:f=0$ which are tame with respect to the subset $\hat{\rho}$, the ideal $B(f,\hat{\rho})$ is the Bourbaki ideal of the $S$-module $D_{0}(f)$ in the sense of the general theory of Bourbaki ideals, see [21] and our discussion in Section 3 below.
In Section 3 we compare our sequence (1.4) to the usual notion of a Bourbaki sequence for a (graded) $S$-module. Theorem (3.3) is the main result here, giving necessary and sufficient conditions for the sequence (1.4) to be exact. The proof of this result uses a key result due to J. Herzog, S. Kumashiro and D. Stamate in [21]. As we shall see, being tame with respect to $\hat{\rho}$ depends on the choice of $\hat{\rho}$, see Example 3.5.
In Section 4 we show that for tame hypersurfaces, many properties of the $S$-module $D_{0}(f)$ which hold for $n=2$ extend to the case $n>3$, see Theorems 4.1 and 4.3, but fail as soon as we omit this property, see Remark 4.2 and Remark 4.4.
We end the paper with a number of natural questions concerning the module of Jacobian derivations, which might foster new research in this area.
2. Basic facts to Jacobian syzygies of curves
When $n=2$, we use the simpler notations $x_{0}=x,\ x_{1}=y$ and $x_{2}=z$. According to Hilbert Syzygy Theorem, the general form of the minimal resolution for the Milnor algebra $M(f)$ of a curve $C:f=0$ that is assumed to be not free, is the following
(2.1) $0\to\oplus_{i=1}^{m-2}S(-e_{i})\to\oplus_{j=1}^{m}S(1-d-d_{j})\to S^{3}(1-d)\to S,$
with $e_{1}\leq\ldots\leq e_{m-2}$ and $d_{1}\leq\ldots\leq d_{m}$. The last morphism in (2.1) is given by
$(a,b,c)\mapsto af_{x}+bf_{y}+cf_{z},$
and has as kernel precisely the graded $S$-module of Jacobian syzygies $D_{0}(f)$ shifted by $(1-d)$. It follows from [20, Lemma 1.1] that one has
(2.2) $e_{j}=d+d_{j+2}-1+\epsilon_{j},$
for $j=1,\ldots,m-2$ and some integers $\epsilon_{j}\geq 1$. With this notation, the resolution in (2.1) yields the following minimal resolution for the graded $S$-module $D_{0}(f)$:
(2.3) $0\to F_{1}=\oplus_{i=1}^{m-2}S(-e^{\prime}_{i})\to F_{0}=\oplus_{j=1}^{m}S(-d_{j}),$
where $e^{\prime}_{j}=d_{j+2}+\epsilon_{j}$, for $j=1,\ldots,m-2$. Therefore, the integers $\epsilon_{j}$ gives the shifts of the degrees of the generators when we pass from $F_{0}$ to $F_{1}$ in the minimal resolution of $D_{0}(f)$. Using [20, Formula (13)], it follows that one has
(2.4) $d_{1}+d_{2}=d-1+\sum_{i=1}^{m-2}\epsilon_{j}\geq d-1+m-2,$
which gives in particular the following bounds on the number $m$ of a minimal generating set
(2.5) $2\leq m\leq d_{1}+d_{2}-d+3,$
see [16, Proposition 2.1] and [19]. On the other hand, for a reduced degree $d$ curve $C$, one has
(2.6) $m\leq d+1,$
see [16, Proposition 2.1], while for the case $C$ a line arrangement, we have the slightly stronger inequality
(2.7) $m\leq d-1,$
see [12, Corollary 1.3], It is known that, for a reduced degree $d$ curve $C$, one has
(2.8) $d_{m}\leq 2d-4,$
see [5, Corollary 11], and for the case $C$ a line arrangement, one has the much stronger inequality
(2.9) $d_{m}\leq d-2,$
see [26, Corollary 3.5].
Next we collect some characterizations of free curves. The first one is very easy to check, as soon as we have the first two exponents of $C$.
(C1)     A free curve $C$ is characterized by the condition $d_{1}+d_{2}=d-1$,
and hence in this case $r=d_{1}<d/2$. For a proof, see [28]. Recall that the total Tjurina number $\tau(C)$ of a reduced plane curve $C:f=0$ is just the degree $\deg\Sigma$ of the singular subscheme, or equivalently the sum of the Tjurina numbers of all the singularities of $C$. The main result in [17], restated below in Theorem 2.1, gives an upper bound for the total Tjurina number $\tau(C)$ of $C$ as a function of $d$ and $r=d_{1}$. Moreover, we know that when $r<d/2$ the upper bound is obtained exactly for a free curve, see Corollary 2.2 below. Moreover, it was shown in [10] that for any $r<d/2$ there is a free curve with exponents $d_{1}=r$ and $d_{2}=d-1-r$.
Theorem 2.1.
For positive integers $d$ and $r$, set $\tau(d,r)_{max}=(d-1)^{2}-r(d-r-1).$ Then, if $C:f=0$ is a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r=d_{1}={\rm mdr}(f)$, one has
$\tau(C)\leq\tau(d,r)_{max}.$
For a proof, see [17, Theorem 3.2] as well as [18] for an alternative approach.
At the end of the proof of Theorem 2.1, in [17], the authors state the following very interesting consequence (of the proof, not of the statement) of Theorem 2.1.
Corollary 2.2.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r=d_{1}={\rm mdr}(f)$. One has
$\tau(C)=\tau(d,r)_{max}$
if and only if $C:f=0$ is a free curve, and then $r<d/2$.
In the paper [8], there is an alternative proof of Corollary 2.2. Moreover, a plane curve $C$ is nearly free, which can be defined by the property $m=3$, $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$ if and only if a similar property holds. Namely, one has the following result, see [8].
Corollary 2.3.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$. One has
$\tau(C)=\tau(d,r)_{max}-1$
if and only if $C:f=0$ is a nearly free curve, and then $r\leq d/2$.
Next we state a characterization of POG curves, which is very easy to check, as soon as we have the first two exponents of $C$.
(C2)     A non free curve $C$ satisfies $d_{1}+d_{2}\geq d$. Moreover, the equality holds if and only if $C$ is a POG curve,
see [15, Theorem 2.3].
The exact sequence (1.4) in the case $n=2$ goes back to [17], and was gradually developed in [8, 13, 14, 22]. The main fact is the following result.
(C3)     Any curve $C:f=0$ in $\mathbb{P}^{2}$ is tame with respect to $\hat{\rho}=(\rho_{1})$, a syzygy of minimal degree in $D_{0}(f)$,
see [8, Lemma 3.2 and the exact sequence (3.3)].
3. Tame hypersurfaces vs. Bourbaki exact sequences
The tameness property plays a key role in this section, so we start with a discussion of this property. First we recall the notion of Bourbaki sequence for a module. Let $R$ be a Noetherian ring and $M$ an $R$-module, where both $R$ and $M$ are not supposed to be graded. Then a Bourbaki sequence of $M$ means a short exact sequence
(3.1) $0\to F\to M\to I\to 0$
of $R$-modules, where $F$ is a free $R$-module and $I$ is an ideal of $R$. It is known that a Bourbaki sequence of $M$ always exists if $R$ is a normal domain and $M$ is a finitely generated torsionfree $R$-module, see [2, Chapter VII, §4, 9. Theorem 6]. Bourbaki sequences were considered also for graded modules. For instance, [21, Theorem 2.1 and Lemma 2.2 ] yields the following result.
Theorem 3.1.
Let $S$ be the polynomial ring $\mathbb{C}[x_{0},\ldots,x_{n}]$ with $n\geq 2$ as above and $M$ a finitely generated reflexive graded $S$-module of rank $r>1$. Then there is a graded Bourbaki sequence
$0\to S(-D)^{r-1}\to M\to I(s)\to 0,$
where $D$ is the maximal degree of a generator of $M$ in a minimal set of generators for $M$, $s$ is an integer and $I$ a graded ideal in $S$ with $\operatorname{grade}(I)=2$.
Note that the sequence (1.4) is not a Bourbaki sequence for $D_{0}(f)$ as in Theorem 3.1, not even when $n=\operatorname{rank}D_{0}(f)=2$, since in this case we usually choose $\hat{\rho}=(\rho_{1})$, with $\rho_{1}$ a minimal degree generator for $D_{0}(f)$. When $n>2$, our sequence (1.4) relates in a more precise way the numerical invariants of the hypersurface $V:f=0$, assumed to be tame with respect to $\hat{\rho}$, to the degrees of the of the syzygies in $\hat{\rho}$, see for instance Theorems 4.1, 4.3 and 4.11 below. However, Theorem 3.1 clearly implies the following result.
Corollary 3.2.
Any reduced hypersurface $V:f=0$ in $\mathbb{P}^{n}$ is tame for some subset $\hat{\rho}$ if the minimal set of generators for $D_{0}(f)$ satisfies
$d_{1}=d_{2}=\ldots=d_{m}.$
In general, the sequence (1.4) is exact at the first term, that is $u$ is injective exactly when the $(n-1)\times(n+1)$-matrix $M(\hat{\rho})^{\prime}$ which has as rows the components of the syzygies in $\hat{\rho}$ satisfies
(3.2) $\operatorname{rank}M(\hat{\rho})^{\prime}=n-1,$
where the rank is computed say over the field of fractions $K$ of $S$. To check whether the sequence (1.4) is exact at the second term, that is
$\operatorname{im}u=\ker v,$
it is more subtle. We have the following result.
Theorem 3.3.
Let $I(\hat{\rho})$ be the ideal in $S$ generated by all the $(n-1)$-minors in the matrix $M(\hat{\rho})^{\prime}$. Then the sequence (1.4) is exact if and only if
$\operatorname{height}(I(\hat{\rho})\geq 2.$
Proof.
The proof uses the key result which appear in [21, Theorem 1.2], and covers the ungraded situation described in (3.1). To apply this result, we recall the following facts. First the polynomial ring $S$ is clearly a normal domain of dimension $n+1\geq 2$, since we have assumed in this paper that $n\geq 2$. The $S$-module $D_{0}f)$ is finitely generated and torsion free, and has rank $n\geq 2$. Moreover, $D_{0}(f)$ is reflexive and sits in the exact sequence of $S$-modules
$0\to D_{0}(f)\to S^{n+1}\to J_{f}(d-1)\to 0,$
which satisfies the hypotheses of [21, Theorem 1.2 (a)]. Then [21, Theorem 1.2 (a)] tells us that $\operatorname{height}(I(\hat{\rho})\geq 2$ if and only if the following sequence of (ungraded) $S$-modules
[…]
is exact and $I^{\prime}=\operatorname{coker}(u)=D_{0}(f)/\operatorname{im}(u)$ is an ideal in $S$. Since the grading plays no role here, we may replace $S(-d_{1})\oplus\ldots\oplus S(-d_{n-1})$ by $S^{n-1}$ to have the formulation exactly as in [21, Theorem 1.2 (a)]. In fact, the injectivity of $u$ is clear, since the condition (3.2) fails if and only if $I(\hat{\rho})=0$, which contradicts the inequality $\operatorname{height}(I(\hat{\rho})\geq 2$, since $\operatorname{height}(0)=0$. If the equality $\operatorname{im}u=\ker v$ fails, then any element $\rho\in\ker v$ such that $\rho\notin\operatorname{im}v$ gives rise to a torsion element in $I^{\prime}$, which is a contradiction with the fact that $I^{\prime}$ is an ideal in $S$. This implies by [21, Theorem 1.2 (a)] that in this case $\operatorname{height}(I(\hat{\rho}))<2$. If the equality $\operatorname{im}u=\ker v$ holds, then the exact sequence (1.4) may be identified to the sequence (3.3), and we conclude, using again [21, Theorem 1.2 (a)], that in this case $\operatorname{height}(I(\hat{\rho})\geq 2$. This completes our proof, since the sequence (1.4) is exact at the third term by the definition of the Bourbaki ideal.
Remark 3.4.
Recall that for any ideal $I\subset S$, one has
$\operatorname{height}(I)=\min_{I\subset\mathfrak{p}}\operatorname{height}(\mathfrak{p})=\min_{I\subset\mathfrak{p}}\operatorname{codim}V(\mathfrak{p}),$
see for instance [4, Theorem A.16], where the minimum is over all prime ideal $\mathfrak{p}$ containing $I$, in other words over all the irreducible varieties $V(\mathfrak{p})$ contained in the scheme $V(I)$ defined by $I$ in $\mathbb{C}^{n+1}$. On the other hand, we clearly have
$\operatorname{codim}V(I)=\min_{I\subset\mathfrak{p}}\operatorname{codim}V(\mathfrak{p}).$
Hence the condition $\operatorname{height}(I(\hat{\rho}))\geq 2$ in Theorem 3.3 is equivalent to
$\operatorname{codim}V(I(\hat{\rho}))\geq 2.$
In other words, $V(I)$ contains no codimension 1 component, that is the set of all the $(n-1)$-minors in the matrix $M(\hat{\rho})^{\prime}$ are not divisible by any polynomial in $S$ of degree $>0$. Similarly, since in the polynomial ring $S$ we have
$\operatorname{grade}(I)=\operatorname{depth}(I)=\operatorname{height}(I)$
for any ideal $I$, it follows that the condition $\operatorname{grade}(I)=2$ in Theorem 3.1 is equivalent to
$\operatorname{codim}I=2.$
Now it turns out that for $n=2$, any reduced plane curve is tame, see the property (C3) in Section 2. For $n=3$, the first condition (3.2) holds for any surface and any choice $\hat{\rho}=(\rho_{1},\rho_{2})$ with $\rho_{1}$ of minimal degree and $\rho_{2}\notin S\cdot\rho_{1}$. For $n\geq 4$, the tameness condition is even more subtle. Here is an example illustrating this property in the case $n=3$, see also Remark 4.4 and Example 4.8 below. We use in this paper the simpler notation $x_{0}=x,\ x_{1}=y,\ x_{2}=z$ and $x_{3}=w$ when $n=3$.
Example 3.5.
The following example is related in particular to Corollary 3.2. Consider the surface
$V:f=(x^{2}-zw)(x^{2}+y^{2}+z^{2})=0,$
which is a union of two quadratic cones in $\mathbb{P}^{3}$. Then the computer algebra software SINGULAR [6] gives $m=7$ and the exponents $(2,2,2,2,2,2,2)$. More precisely, we have
$\rho_{1}=(0,-z^{2},yz,-yw),\ \rho_{2}=(z^{2},0,-xz,2xz+xw),\ \rho_{3}=(yw,-2xz-xw,2xy,0),$
$\rho_{4}=(yz,-xz,0,2xy),\rho_{5}=(-xy,2x^{2}+y^{2}+z^{2},0,-2yw),$
$\rho_{6}=(-2xz+xw,2yz+yw,4x^{2}+2z^{2}+zw,-6zw-3w^{2})$
and
$\rho_{7}=(4xz-xw,-yw,-4x^{2}-zw,8x^{2}+3w^{2}).$
It follows that $z\rho_{4}=x\rho_{1}+y\rho_{2}$. Hence, if we choose the subset $\hat{\rho}=(\rho_{1},\rho_{2})$, then $V$ is not tame with respect to $\hat{\rho}$, since $v(\rho_{4})=0$ and clearly $\rho_{4}\notin\operatorname{im}u$. However, if we take the subset $\hat{\rho}=(\rho_{1},\rho_{6})$, a direct computation with SINGULAR shows that $\operatorname{height}(I(\hat{\rho}))=2$, and hence $V$ is tame with respect to this $\hat{\rho}$ by Theorem 3.3.
4. Results related to Jacobian syzygies of hypersurfaces
It is clear that any free hypersurface is tame, for any choice of the first $n-1$ generators $\hat{\rho}$. The following result can be regarded as an extension of property (C1) from previous section to dimensions $n>2$.
Theorem 4.1.
Let $V:f=0$ be a tame hypersurface with respect to the subset $\hat{\rho}=(\rho_{1},\ldots,\rho_{n-1})$ and having the (unordered) set of exponents $(d_{1},\ldots,d_{m})$ for some $m\geq n$. Then the following hold.
  1. (1)
    The Bourbaki ideal $B(f,\hat{\rho})\subset S$ is saturated and $\dim S/B(f,\hat{\rho})\leq n-1$. Moreover, the minimal number of generators of $B(f,\hat{\rho})$ is $m-n+1$.
  2. (2)
    The hypersurface $V:f=0$ is free if and only if $B(f,\hat{\rho})=S$.
  3. (3)
    The hypersurface $V:f=0$ is an $(n+1)$-syzygy hypersurface if and only if the ideal $B(f,\hat{\rho})$ is a codimension 2 complete intersection.
  4. (4)
    For any $j\geq n$ one has
    $d_{1}+\ldots+d_{n-1}+d_{j}\geq d-1,$
    and the equality holds for some $j\geq n$ if and only if the tame hypersurface $V$ is free.
Proof.
To prove (1), let $\delta=d_{1}+\ldots+d_{n-1}-(d-1).$ Then the exact sequence (1.4) can be written in the form
[…]
By taking the associated coherent sheaves, we get the following exact sequence
[…]
Looking at the corresponding long exact sequence of cohomology groups, and using the fact that
$H^{0}(\mathbb{P}^{n},\tilde{B}(f,\hat{\rho}))=B(f,\hat{\rho})^{s},$
the saturation of the ideal $B(f,\hat{\rho})$, and
$H^{1}(\mathbb{P}^{n},{\mathcal{O}}(u))=0$
for any $n\geq 2$ and any integer $u$, we get the exact sequence
[…]
By comparing (4.1) and (4.2), we see that $B(f,\hat{\rho})=B(f,\hat{\rho})^{s}$, hence the Bourbaki ideal is saturated. To prove that $\dim S/B(f,\hat{\rho})\leq n-1$ one can use [21, Lemma 2.2], since the $S$-module $D_{0}(f)$ is reflexive by definition. The claim about the minimal number of generators follows from the exact sequence (4.1), which shows that any system of generators for $B(f,\hat{\rho})$ lifts to a set of syzygies which together with the subset $\hat{\rho}$ generate $D_{0}(f).$ This remark completes the proof of (1) and also proves (3). Indeed, an ideal $I\subset S$ generated by two homogeneous polynomial such that $\dim S/I\leq n-1$ is a complete intersection of dimension $n-1$. To prove (2) and (4), note that if $d_{1}+\ldots+d_{n-1}+d_{j}<d-1$, then $e=\delta+d_{j}<0$. By taking the $e$-homogeneous component of the sequence (4.1), we get $v(\rho_{j})=0$ since $S_{e}=0$. By tameness, this implies that $\rho_{j}\in\operatorname{im}u$, a contradiction with the minimality of the generating set $\rho_{k}$’s. If the equality holds, then $e=0$ and $v(\rho_{j})$ is a non-zero constant, which implies that $B(f,\hat{\rho})=S$. A well-known result by K. Saito [25], see for instance [9, Theorem 8.1] for a formulation in our setting, shows that this happens if and only if $V$ is free. ∎
Remark 4.2.
Without the tameness hypothesis, there are non-free hypersurfaces such that
$d_{1}+\ldots+d_{n}=d-1,$
see Example 4.8 below.
The following result can be regarded as an extension of Corollary 2.2 to dimensions $n>2$. In the case $n=3$ this result is essentially [7, Theorem 1.4].
Theorem 4.3.
Let $V:f=0$ be a free hypersurface in $\mathbb{P}^{n}$ with exponents $(d_{1},\ldots,d_{n})$ and $d=\deg f$. Then the singular subscheme $\Sigma$ has dimension $n-2$ and its degree is given by
$\deg\Sigma=(d-1)^{2}-s_{2},$
where $s_{2}=\sum_{1\leq i<j\leq n}d_{i}d_{j}$. Moreover, let $V^{\prime}:f^{\prime}=0$ in $\mathbb{P}^{n}$ be a tame hypersurface with respect to the subset $\hat{\rho}^{\prime}=(\rho_{1}^{\prime},\ldots,\rho_{n-1}^{\prime})$, of degree $d^{\prime}$ and with exponents $(d_{1}^{\prime},\ldots,d_{m}^{\prime})$ for some $m\geq n$. If the singular subscheme $\Sigma^{\prime}$ of $V^{\prime}$ has dimension $n-2$, then the degree of the singular subscheme $\Sigma^{\prime}$ satisfies the inequality
$\deg\Sigma^{\prime}\leq(d^{\prime}-1)^{2}-\sum_{1\leq i<j\leq n}d_{i}^{\prime}d_{j}^{\prime}$
where in the last sum $d_{n}^{\prime}$ is replaced by
$d_{n}^{\prime\prime}=d^{\prime}-1-\sum_{j=1}^{n-1}d_{j}^{\prime}.$
The equality holds here if and only if $V^{\prime}$ is free and hence in particular $m=n$.
Proof.
Since $V$ is a free hypersurface in $\mathbb{P}^{n}$ with exponents $(d_{1},\ldots,d_{n})$, it follows that
(4.3) $D_{0}(f)=\oplus_{j=1}^{n}S(-d_{j})$
and hence, for any large integer $k$ we have
$\dim D_{0}(f)_{k}=\sum_{j=1}^{n}\binom{k+n-d_{j}}{n}.$
The obvious exact sequence
(4.4) $0\to D_{0}(f)\to S^{n+1}\to S(d-1)\to M(f)(d-1)\to 0$
implies that
$\dim D_{0}(f)_{k}=(n+1)\binom{k+n}{n}-\binom{k+n+d-1}{n}+\dim M(f)_{k+d-1}.$
Therefore we get
(4.5) $\dim M(f)_{k+d-1}=\sum_{j=1}^{n}\binom{k+n-d_{j}}{n}-(n+1)\binom{k+n}{n}+\binom{k+n+d-1}{n}.$
Define polynomials $A_{j}(t)\in\mathbb{Z}[t]$ for $j=1,\ldots,n$ such that
$n!\binom{k+n}{n}=k^{n}+A_{1}(n)k^{n-1}+A_{2}(n)k^{n-2}+\ldots,$
where $\ldots$ denotes terms in $k$ with exponent $<n-2$. With this notation, one has
$n!\binom{k+n-d_{j}}{n}=$
$k^{n}+(A_{1}(n)-nd_{j})k^{n-1}+(A_{2}(n)-(n-1)d_{j}A_{1}(n)+\binom{n}{2}d_{j}^{2})k^{n-2}+\ldots$
and
$n!\binom{k+n+d-1}{n}=$
$k^{n}+(A_{1}(n)-n+nd)k^{n-1}+(A_{2}(n)+(n-1)(d-1)A_{1}(n)+\binom{n}{2}(d-1)^{2})k^{n-2}+\ldots.$
Using the equality (4.5) it follows that
$n!\dim M(f)_{k+d-1}=n(d-1-\sum_{j=1}^{n}d_{j})k^{n-1}+$
$+\left((n-1)(d-1-\sum_{j=1}^{n}d_{j})A_{1}(n))+\binom{n}{2}((d-1)^{2}+\sum_{j=1}^{n}d_{j}^{2})\right)k^{n-2}+\ldots.$
Comparing this formula with the equality (1.2), we see that the coefficient of $k^{n-1}$ has to vanish, and hence we get
$d-1=\sum_{j=1}^{n}d_{j},$
as we have already seen in Theorem 4.1. It follows that the coefficient of $k^{n-2}$ is strictly positive and hence $s=\dim\Sigma=n-2$ and
$\deg\Sigma=\frac{(n-2)!}{n!}\binom{n}{2}((d-1)^{2}+\sum_{j=1}^{n}d_{j}^{2})=\frac{(\sum_{j=1}^{n}d_{j})^{2}+\sum_{j=1}^{n}d_{j}^{2}}{2}=s_{1}^{2}-s_{2},$
where $s_{1}=\sum_{j=1}^{n}d_{j}$ and $s_{2}=\sum_{1\leq i<j\leq n}d_{i}d_{j}$. Finally, consider a tame hypersurface $V^{\prime}:f^{\prime}=0$. Then the sequence of graded $S$-modules (1.4) is exact. This implies that
$\dim D_{0}(f)_{k}\leq\sum_{j=1}^{n-1}\binom{k+n-d_{j}^{\prime}}{n}+\binom{k+n-d^{\prime}+1+\sum_{j=1}^{n-1}d_{j}^{\prime}}{n},$
with equality if and only if the corresponding Bourbaki ideal $B(f^{\prime},\hat{\rho}^{\prime})$ is the polynomial ring $S$ by Theorem 4.1. Using as above the exact sequence (4.4) it follows that
$\dim M(f^{\prime})_{k+d^{\prime}-1}\leq\sum_{j=1}^{n-1}\binom{k+n-d_{j}^{\prime}}{n}+\binom{k+n-d^{\prime}+1+\sum_{j=1}^{n-1}d_{j}^{\prime}}{n}-$
$-(n+1)\binom{k+n}{n}+\binom{k+n+d^{\prime}-1}{n}.$
Using the above notation, we see that
$n!\dim M(f^{\prime})_{k+d^{\prime}-1}\leq(d^{\prime}-1)^{2}-\sum_{1\leq i<j\leq n}d_{i}^{\prime}d_{j}^{\prime}$
where $d_{n}^{\prime}$ is replaced by $d^{\prime}-1-\sum_{j=1}^{n-1}d_{j}^{\prime}.$
Remark 4.4.
1) Without the tameness hypothesis with respect to $\hat{\rho}^{\prime}$, there are non-free hypersurfaces $V^{\prime}$ such that the inequality for $\deg\Sigma^{\prime}$ is an equality, see Example 4.8 below. Moreover, the inequality for $\deg\Sigma^{\prime}$ holds even when $d_{n}^{\prime\prime}\leq 0$, see Examples 3.5 and 4.6.
2) Note that there are tame hypersurfaces $V$ such that the corresponding singular subscheme satisfies $\dim\Sigma<n-2$. For instance, the surface in $\mathbb{P}^{3}$ given by
$V:f=xyz-w^{3}=0$
has isolated singularities. Them $m=5$ and the generating syzygies are
$\rho_{1}=(0,y,-z,0),\rho_{2}=(x,0,-z,0),\ \rho_{3}=(3w^{2},0,0,yz),$
$\rho_{4}=(0,3w^{2},0,xz)\text{ and }\rho_{5}=(0,0,3w^{2},xy).$
Then it is easy to see that $V$ is tame with respect to $\hat{\rho}=(\rho_{1},\rho_{2})$. Indeed, the matrix $M(\hat{\rho})$ has rank 2 as required in (3.2). To show that $\ker v=\operatorname{im}u$ also holds, we proceed as follows. We can take $\rho=A\rho_{3}+B\rho_{4}+C\rho_{5}$ such that $A,B,C\in S$ and $v(\rho)=0$. A direct computation shows that this means
$Ayz+Bxz+Cxy=0.$
This implies $A=xA_{1}$, $B=yB_{1}$ and $C=zC_{1}$ with $A_{1},B_{1},C_{1}\in S$. It follows that $A_{1}+B_{1}+C_{1}=0$. Therefore one has
$\rho=A_{1}(x\rho_{3}-z\rho_{5})+B_{1}(y\rho_{4}-z\rho_{5}).$
Since both $(x\rho_{3}-z\rho_{5})$ and $(y\rho_{4}-z\rho_{5})$ are clearly in $\operatorname{im}u$, it follows that $\rho\in\operatorname{im}u$, and this proves our claim.
The following result is a partial extension of the characterization (C2) of POG curves given in the previous section to higher dimensions. For the partial extension of the converse implication see Theorem 4.11 below.
Theorem 4.5.
Let $V:f=0$ be a strictly POG hypersurface with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$ with exponents $(d_{1},\ldots,d_{n},d_{n+1})$. Then $d=\deg f$ is given by
$d=d_{1}+\ldots+d_{n}$
and, if $\dim\Sigma=n-2$, then
$\deg\Sigma=\sum_{j=1}^{n}d_{j}^{2}+\sum_{1\leq i<j\leq n}d_{i}d_{j}-\sum_{j=1}^{n+1}d_{j}.$
In particular $v(\rho_{n})\neq 0$ and $\deg v(\rho_{n})=1$.
The equality $d=d_{1}+\ldots+d_{n}$ for a POG hyperplane arrangement was obtained already in [1, Proposition 4.1].
Proof.
To prove the equality $d=d_{1}+\ldots+d_{n}$, we use the same aproach as in the proof of Theorem 4.3. The equality (4.3) has to be replaced by the exact sequence
$0\to S(-d_{n+1}-1)\to\oplus_{j=1}^{n+1}S(-d_{j})\to D_{0}(f)\to 0,$
which follows from the definition of a POG hypersurface. This yields the following
$\dim D_{0}(f)_{k}=\sum_{j=1}^{n+1}\binom{k+n-d_{j}}{n}-\binom{k+n-d_{n+1}-1}{n}.$
Next, using the exact sequence (4.4), we get the following equality which replaces (4.5).
$\dim M(f)_{k+d-1}=\sum_{j=1}^{n+1}\binom{k+n-d_{j}}{n}-\binom{k+n-d_{n+1}-1}{n}-$
$-(n+1)\binom{k+n}{n}+\binom{k+n+d-1}{n}.$
The right hand side of this equality is shown to be a polynomial in $k$ of degree $n-1$ with leading coefficient
$c_{n-1}=d-(d_{1}+\ldots+d_{n}).$
By comparing to (1.2), we see that this leading coefficient is in fact zero. The coefficient of $k^{n-2}$ is given by
$c_{n-2}=\frac{1}{n!}\binom{n}{2}(\sum_{j=1}^{n}d_{j}^{2}+d^{2}-2d-2d_{n+1}).$
If $\dim\Sigma=n-2$, then we get
$\deg\Sigma=(n-2)!c_{n-2}=\sum_{j=1}^{n}d_{j}^{2}+\sum_{1\leq i<j\leq n}d_{i}d_{j}-\sum_{j=1}^{n+1}d_{j}.$
If $\dim\Sigma<n-2$, then
$\sum_{j=1}^{n}d_{j}^{2}+\sum_{1\leq i<j\leq n}d_{i}d_{j}-\sum_{j=1}^{n+1}d_{j}=0,$
and hence in this case $d_{n+1}$ is determined by the $d_{j}$’s with $j\leq n$. It remains to show that $v(\rho_{n})\neq 0$. Note that $v(\rho_{n})=0$ means that there is a non-trivial second order syzygy
$\delta=a_{0}E+a_{1}\rho_{1}+\ldots+a_{n}\rho_{n}=0$
involving the Euler derivation
$E=(x_{0},x_{1},\ldots,x_{n})$
and the syzygies $\rho_{j}\in D_{0}(f)$. By applying $\delta$ to $f$, we see that $a_{0}=0$, and hence we get a non-trivial second order syzygy
$\delta=a_{0}E+a_{1}\rho_{1}+\ldots+a_{n}\rho_{n}=0$
involving the syzygies $\rho_{j}\in D_{0}(f)$ for $j<n+1$, and hence not a multiple of the syzygy in (1.3). This contradiction proves our final claim. ∎
Example 4.6.
Consider the surface
$V:f=(x^{3}-yzw)^{3}+(w^{3}-xyz)^{3}=0,$
which is a union of 3 cubic surfaces in $\mathbb{P}^{3}$. Then the computer algebra software SINGULAR [6] gives $m=8$ and the exponents $(1,4,4,7,7,8,8,8)$. In particular we have
$d_{1}+d_{2}+d_{3}=1+4+4=9=\deg V,$
but $V$ is far from being a POG surface. The Hilbert polynomial for $M(f)$ is $11938t-$ 119, which shows that $\dim\Sigma=1$. The software SINGULAR [6] gives the following first 3 generating syzygies.
$\rho_{1}=(0,y,-z,0),$
$\rho_{2}=(xyz^{2}+3zw^{3},9x^{2}w^{2},-yz^{3},3x^{3}z+yz^{2}w)$
and
$\rho_{3}=(xy^{2}z+3yw^{3},0,-y^{2}z^{2}+9x^{2}w^{2},3x^{3}y+y^{2}zw).$
For both pairs $\hat{\rho}=(\rho_{1},\rho 2)$ and $\hat{\rho}=(\rho_{1},\rho_{3})$ we get $\operatorname{height}(I(\hat{\rho}))=1<2$, hence Theorem 3.3 tells us that $V$ is not tame with respect to these two choices for $\hat{\rho}$. However, consider the last generating syzygy listed by SINGULAR, namely $\rho_{8}=(A,B,C,D)$, where
$A=x^{6}yz-3x^{2}y^{2}z^{2}w^{2}+y^{3}z^{3}w^{2}+6x^{2}yzw^{4}+6xyzw^{5}-3w^{8},\ B=0$
$C=-2x^{2}y^{2}z^{3}w+18x^{4}zw^{3}\text{ and }D=3x^{8}-x^{2}y^{3}z^{3}+5x^{2}y^{2}z^{2}w^{2}+2xy^{2}z^{2}w^{3}-yzw^{6}).$
If we take now $\hat{\rho}=(\rho_{1},\rho_{8})$, a direct computation by the computer algebra software SINGULAR shows that $\operatorname{height}(I(\hat{\rho})=2$, and hence $V$ is tame with respect to this $\hat{\rho}$ by Theorem 3.3.
The next result shows that strict POG hypersurfaces are intimately related to tame POG hypersurfaces.
Proposition 4.7.
Let $V:f=0$ be a POG hypersurface with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$. Then the hypersurface $V:f=0$ is strictly POG with respect to $\overline{\rho}$ and $\ell$ does not divide $b_{n}$ if and only if $V:f=0$ is tame with respect to the subset $\hat{\rho}=(\rho_{1},\ldots,\rho_{n-1})$.
Proof.
Suppose first that $V$ is strictly POG and $\ell$ does not divide $b_{n}$ in (1.3). If condition (3.2) fails, then this is equivalent to the existence of a non trivial relation of the form
$c_{1}\rho_{1}+\ldots+c_{n-1}\rho_{n-1}=0.$
Such a relation cannot be a multiple of the relation (1.3), since $\ell\neq 0$. Hence the condition (3.2) holds. Let now
$\rho=\sum_{j=1}^{n+1}c_{j}\rho_{j}$
be a syzygy not in $\operatorname{im}u$, and hence at least one of $c_{n}$ and $c_{n+1}$ is non-zero. We can assume moreover that $\ell$ does not divide $c_{n+1}$ in view of (1.3). Over the field of fractions $K$ of the polynomial ring $S$, we know as above that $\rho_{1},\ldots,\rho_{n-1}$ are linearly independent. If $\rho_{n}$ is in the $K$-vector space spanned by $\rho_{1},\ldots,\rho_{n-1}$, then the $K$-vector space spanned by $\rho_{1},\ldots,\rho_{n+1}$ would be of dimension $n-1$, a contradiction, since this dimension equals the rank of $D_{0}(f)$ and is equal to $n$. It follows that $v(\rho_{n})\neq 0$. Assume now that $v(\rho)=0$, hence
$v(\rho)=c_{n}v(\rho_{n})+c_{n+1}v(\rho_{n+1})=0.$
By applying $v$ to (1.3) we get
$b_{n}v(\rho_{n})+\ell v(\rho_{n+1})=0.$
Using $v(\rho_{n})\neq 0$, it follows that
$c_{n}\ell=c_{n+1}b_{n}.$
This is a contradiction, since $\ell$ divides neither $b_{n}$ nor $c_{n+1}$. Conversely, assume that $V:f=0$ is tame. If $\ell=0$, then by taking $\rho=\rho_{n}$ we see that the equality $\ker v=\operatorname{im}u$ does not hold. Indeed, $\rho_{n}\notin\operatorname{im}u$, since we have a minimal set of generators. On the other hand, the equality (1.3) with $\ell=0$ implies that the equality $\ker v=\operatorname{im}u$ fails.
Example 4.8.
Consider the surface
$V:f=(x^{2}+y^{2}-zw)^{2}-2w^{4}=0.$
Then the computer algebra software SINGULAR [6] gives the following syzygies
$\rho_{1}=(0,w,2y,0),\ \rho_{2}=(-y,x,0,0),\rho_{3}=(w,0,2x,0)$
and
$\rho_{4}=(xz,yz,2z^{2}-8w^{2},2x^{2}+2y^{2}-2zw).$
There is just one generator for the second order syzygies, namely the relation
$x\rho_{1}-w\rho_{2}-y\rho_{3}=0.$
It follows that $V:f=0$ is, for this ordering of the syzygies $\rho_{j}$’s, a non strictly POG surface. This shows in particular that the hypothesis strictly POG is necessary in Theorem 4.5. Indeed, without this hypothesis, the first exact sequence in the proof of Theorem 4.5 might not be the minimal resolution of $D_{0}(f)$. Note also that $V$ is not tame with respect to $\hat{\rho}=(\rho_{1},\rho_{2}).$ However, if we set $\rho_{1}^{\prime}=\rho_{4}$ and $\rho_{4}^{\prime}=\rho_{1}$, and consider the syzygies in the order $\rho_{1}^{\prime},\rho_{2},\rho_{3},\rho_{4}^{\prime}$, then $V:f=0$ is a strictly POG surface, and even a tame surface by Proposition 4.7, with exponents $(2,1,1,1)$. Using Theorem 4.5, we see that
$\deg\Sigma=4+1+1+2+2+1-(2+1+1+1)=6.$
This shows that we have equality in the last inequality in Theorem 4.3, where $d_{3}^{\prime\prime}=1$ if we use the 3 syzygies of degree 1, even if $V$ is not free. This is due to the fact that $V$ is not tame with respect to $\hat{\rho}=(\rho_{1},\rho_{2})$, but only with respect to $\hat{\rho}^{\prime\prime}=(\rho_{1}^{\prime},\rho_{2})$. If we use $d_{1}^{\prime}=2$ and $d_{2}^{\prime}=1$, then $d_{3}^{\prime\prime}=0$ and hence the inequality given by Theorem 4.3 becomes
$6=\deg\Sigma<9-1\times 2=7,$
in accord with the fact that $V$ is not free.
The proof of Proposition 4.7 and Theorem 4.5 yield the following.
Corollary 4.9.
If $V:f=0$ is a non strictly POG with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$, then one of the following two cases occurs.
  1. (1)
    If $b_{n}\neq 0$, then $B(f,\hat{\rho})$ is a principal ideal generated by the polynomial $v(\rho_{n+1})$.
  2. (2)
    If $b_{n}=0$, then $B(f,\hat{\rho})=0$.
If $V:f=0$ is a strictly POG with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$, then one of the following two cases occurs.
  1. (i)
    The Bourbaki ideal $B(f,\hat{\rho})$ defines a codimension 2 complete intersection of multi-degree $(1,d_{n+1}-d_{n}+1)$ if $V$ is tame with respect to the subset $\hat{\rho}$.
  2. (ii)
    The Bourbaki ideal $B(f,\hat{\rho})$ defines a hyperplane if $V$ is not tame with respect to the subset $\hat{\rho}$.
In particular, a POG hypersurface as above is tame if and only if the corresponding Bourbaki ideal $B(f,\hat{\rho})$ defines a codimension 2 complete intersection of multi-degree
$(1,d_{n+1}-d_{n}+1).$
The last claim in Corollary 4.9 may be regarded as a refinement of the last claim in Theorem 3.1 saying that $\operatorname{grade}(B(f,\hat{\rho}))=2$.
Proof.
Consider first the case of a non strictly POG such that $b_{n}\neq 0$. This implies $v(\rho_{n})=0$. On the other hand, note that $v(\rho_{n+1})\neq 0$, since otherwise we get a second order syzygy
$\delta=a_{0}E+a_{1}\rho_{1}+\ldots+a_{n-1}\rho_{n-1}+a_{n+1}\rho_{n+1}=0$
as in the proof of Theorem 4.5 above. As above, we have $a_{0}=0$ and this leads to a contradiction with the assumption $b_{n}\neq 0$. Consider now the case of a non strictly POG such that $b_{n}=0$. Then the syzygies in $\hat{\rho}$ are linearly dependent over the fields of fractions $K$, and hence the map $v$ is the trivial map in this case.
Assume now that is a strictly POG as in (1.3). Then we know that $\ell_{1}=v(\rho_{n})$ is a non-zero linear form by Theorem 4.5. If $\ell$ does not divide $b_{n}$, then $\ell$ divides $\ell_{1}$, and $v(\rho_{n+1})$ is just $b_{n}$ up to a nonzero factor. Hence $B(f,\hat{\rho})=(\ell,b_{n})$ in this case.
If $\ell$ divides $b_{n}$, then $\ell_{1}$ divides $v(\rho_{n+1})$ and hence $B(f,\hat{\rho})=(\ell_{1})$ in this case.
Example 4.10.
Consider the surface
$V:f=(x+y+z)(x+2y)wz+xyzw=0$
which is a union of two planes and a quadratic surface in $\mathbb{P}^{3}$. Then the computer algebra software SINGULAR [6] shows that $m=4$ and the exponents are $(1,1,2,2)$. The Bourbaki ideal corresponding to the choice $\hat{\rho}=(\rho_{1},\rho_{2})$, the two syzygies of degree 1, is given by
$B(f,\hat{\rho})=(y,2x+z),$
and the minimal resolution of the Jacobian ideal is
$0\to S(-6)\to S(-4)^{2}\oplus S(-5)^{2}\to S(-3)^{4}-->S.$
It follows that $V$ is a tame, nearly free surface.
Theorem 4.11.
Let $V:f=0$ be a tame hypersurface with respect to $\hat{\rho}=(\rho_{1},\ldots,\rho_{n-1})$, which is an $(n+1)$-syzygy hypersurface with exponents $(d_{1},\ldots,d_{n+1})$ such that $d=\deg f$ is given by
$d=d_{1}+\ldots+d_{n}.$
Then $V$ is a strictly POG hypersurface with respect to $(\rho_{1},\ldots,\rho_{n+1})$.
Proof.
It follows from Remark 3.4 that the scheme $Z$ defined by the Bourbaki ideal $B(f,\hat{\rho})$ has codimension 2. Exactly as in the final step of the proof of Theorem 4.5, we see that $\ell_{1}=v(\rho_{n})$ is a non-zero linear form. Since $V$ is an $(n+1)$-syzygy hypersurface, it follows that the second order syzygies of $D_{0}(f)$ are generated by a unique relation
(4.6) $a_{1}\rho_{1}+\ldots+a_{n-1}\rho_{n-1}+a_{n}\rho_{n}+a_{n+1}\rho_{n+1}=0,$
where both $a_{n}$ and $a_{n+1}$ are non-zero polynomials of degree $>0$. This relation implies
$a_{n}v(\rho_{n})+a_{n+1}v(\rho_{n+1})=0.$
Since $B(f,\hat{\rho})=(v(\rho_{n}),v(\rho_{n+1}))$ has codimension 2, if $\deg a_{n+1}>1$, then $a_{n+1}=\ell_{1}b$ and $a_{n}=-bv(\rho_{n+1})$, with $\deg b>0$. This implies that
$a_{1}\rho_{1}+\ldots+a_{n-1}\rho_{n-1}-b(v(\rho_{n+1})\rho_{n}+v(\rho_{n})\rho_{n+1})=0,$
and hence
$v(\rho_{n+1})\rho_{n}+v(\rho_{n})\rho_{n+1}\in\ker v=\operatorname{im}u.$
This contradicts the fact that (4.6) generates the second order syzygies of $D_{0}(f)$, and hence proves our claim. ∎
5. Some questions related to Jacobian syzygies of hypersurfaces
Question 5.1.
Find an upper bound for $m$, the minimal number of generators for the graded $S$-module $D_{0}(f)$ in terms of $d$ and $n$. When $V:f=0$ is a generic hyperplane arrangement in $\mathbb{P}^{n}$ then it follows from [24, Theorem 4.5.3 and Corollary 4.5.4] that one has
$m=\binom{d-1}{n-1}.$
Does the bound
(5.1) $m\leq\binom{d-1}{n-1}$
hold for any arrangement of $d$ hyperplanes in $\mathbb{P}^{n}$ The answer is positive for $n=2$, see (2.7). When $V$ is a degree $d$ hypersurface in $\mathbb{P}^{n}$ having only isolated singularities, the following bound
(5.2) $m\leq\binom{n+1}{2}+\deg(\Sigma)$
seems to work. Note that we have equality for a smooth hypersurface $V:f=0$ in $\mathbb{P}^{n}$. Moreover, when $V$ has only isolated singularities, one has
$\deg(\Sigma)=\tau(V),$
the total Tjurina number of $V$, and the above upper bounds seems to hold, compare to [5, Corollary 11]. Note that the homology group $H_{1}(K)$ which appears in that paper is, up to a shift, the quotient of $D_{0}(f)$ by the $S$-submodule generated by all Koszul type derivations of $f$, and there are clearly $\binom{n+1}{2}$ such derivations, all of them of degree $(d-1)$.
It remains of course the problem to bound $\deg(\Sigma)$ in terms of $n$ and $d$. When $V$ has at most isolated singularities one has
$\deg(\Sigma)=\tau(V)\leq(d-1)^{n}.$
On the other hand, when $V$ is a hyperplane arrangement in $\mathbb{P}^{n}$, and $V^{\prime}$ is the intersection of $V$ with a generic 2-plane $E$ in $\mathbb{P}^{n}$, it follows [31, Remark 2.3] that one has
$\deg(\Sigma)=\tau(V^{\prime})\leq(d-1)^{2}.$
It follows that the two upper bounds (5.1) and (5.2) are quite different in nature.
Example 5.2.
Consider the surface
$V:f=x^{4}+y^{4}+z^{4}-xyzw=0.$
A direct computation using SINGULAR shows that $m=12$ and the exponents are $(3,3,3,3,3,3,4,4,4,4,4,4)$. Note that $V$ has isolated singularities and
$\tau(V)=\deg(\Sigma)=10,$
and hence the inequality
$12\leq 6+10$
suggested in Question 5.1 holds.
All examples we have computed so far suggest that the following question has a positive answer
Question 5.3.
Find an upper bound for the maximal degree $d_{m}$ of a Jacobian syzygy in terms of $d$. Does the bound
$d_{m}\leq n(d-2),$
compare to (2.8), works for any $n$ We know that this bound works when $\dim\Sigma=0$, see [5, Corollary 11].
Concerning the tame hypersurfaces, the following two key questions are related.
Question 5.4.
Does there exist hypersurfaces $V:f=0$ which are not tame for any choice of a subset $\hat{\rho}$ of $n-1$ generators in a minimal set of generators for the graded $S$-module $D_{0}(f)$
Theorem 3.1 says that any hypersurface $V:f=0$ is tame with respect to $\hat{\rho}^{\prime}=(\rho_{1}^{\prime},\ldots,\rho_{n-1}^{\prime})$, where $\deg\rho_{j}^{\prime}=d_{m}$, for any $j=1,\ldots,n-1$. Such a subset $\hat{\rho}^{\prime}$ is not a subset of $n-1$ generators in a minimal set of generators for the graded $S$-module $D_{0}(f)$. If the answer to the Question 5.4 is negative, then for any hypersurface $V$ we can define a new invariant $b(V)$, namely the minimal degree of the subset
$\deg(\hat{\rho})=d_{1}+\ldots+d_{n-1},$
where $\hat{\rho}=(\rho_{1},\ldots,\rho_{n-1})$, $d_{j}=\deg\rho_{j}$ and $\hat{\rho}$ runs through the set of subsets of generators such that $V$ is tame with respect to $\hat{\rho}$.
We have seen in Proposition 4.7 and in Example 4.8 that being a strictly POG hypersurface depends on the choice of the first $(n-1)$ syzygies in $\hat{\rho}$. A similar example is the following.
Example 5.5.
Consider the 3-fold
$V:f=(x^{2}+y^{2}+z^{2})(u^{2}+v^{2})=0,$
which is a union of two planes and a quadratic cone in $\mathbb{P}^{4}$. Then the computer algebra software SINGULAR [6] gives $m=5$ and the exponents $(1,1,1,1,1)$, namely one can take
$\rho_{1}=(0,0,0,-v,u),\ \rho_{2}=(0,-z,y,0,0),\ \rho_{3}=(x,y,z,u,v),$
$\rho_{4}=(-y,x,0,0,0)\text{ and }\rho_{5}=(-z,0,x,0,0).$
The minimal resolution of the Jacobian ideal is
$0\to S(-5)\to S(-4)^{5}\to S(-3)^{5}\to S,$
hence $V$ is a nearly free 3-fold. On the other hand, we have the following second order syzygy
$x\rho_{2}+z\rho_{4}-y\rho_{5}=0.$
It follows that for the choice $\hat{\rho}=(\rho_{2},\rho_{4},\rho_{5})$, the 3-fold $V$ is not a strictly POG hypersurface and the corresponding Bourbaki ideal $B(f,\hat{\rho})$ is trivial. But for the choice $\hat{\rho}^{\prime}=(\rho_{1},\rho_{2},\rho_{3})$, the 3-fold $V$ is a strictly POG hypersurface.
Question 5.6.
1) Does there exist a POG hypersurface $V:f=0$ with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$ which is not strictly POG for any choice of a subset $\hat{\rho}$ of $n-1$ generators in a minimal set of generators for the graded $S$-module $D_{0}(f)$
2) Does there exist a POG hypersurface with respect to $\overline{\rho}=(\rho_{1},\ldots,\rho_{n+1})$ which is not tame with respect for any choice of a subset $\hat{\rho}$ of $n-1$ generators in a minimal set of generators for the graded $S$-module $D_{0}(f)$
3) Same questions as 1) and 2), but restricted to the class of nearly free hypersurfaces.
In relation with the example of a tame hypersuface $V$ with $\dim\Sigma<n-2$ given in Remark 4.4, one may ask the following. A negative answer to question 1) below would simplify the hypotheses of Theorem 4.5 as well.
Question 5.7.
1) Does there exist a POG hypersurface such that $\dim\Sigma<n-2$
2) Does there exist a tame POG hypersurface such that $\dim\Sigma<n-2$